# YOLO-TLA: An Efficient and Lightweight Small Object Detection Model based on YOLOv5

### v1

@@ Abstract
Index Terms:
Object detection, deep learning, attention mechanism, neural network.
I Introduction
@@
Currently, there are two principal methodologies to implementing object detection models: the two-stage and the single-stage methods. The two-stage detection process starts with extracting image features using a convolutional neural network (CNN), followed by the generation of multiple candidate bounding boxes (regions) on the feature map. Each box is then processed by additional convolutional layers for object classification and bounding box refinements. This includes identifying the object class and verifying its presence in the bounding box, and a regression phase to enhance the detection accuracy. The process concludes with non-maximum suppression (NMS), which filters out surplus bounding boxes to identify the most confident and accurate final detections. In contrast, the single-stage detection method directly predicts object class and their locations in the image, foregoing candidate bounding boxes generation. This method is generally faster and less computationally intensive than the two-stage method, making it ideal for scenarios requiring high real-time performance.
@@
Since its introduction, the YOLO series of object detection methods has garnered significant attention in the computer vision community, with several iterations enhancing its performance in both academic and industrial settings [9, 10, 11]. The latest detection model, YOLOv5, proposed by Ultralytics Inc., boasts improved accuracy and a more streamlined network architecture. Despite the advancements in YOLOv5, including its balance of speed and accuracy, it confronts specific challenges in industrial applications [12]. The model often struggles with accurately predicting the positions and classes of small or dense objects, negatively impacting detection performance. Furthermore, deploying these models in embedded systems, which typically have limited resources, necessitates a reduction in the number of parameters and computational demands. Consequently, optimizing the model to be more resource-efficient while preserving accuracy is a critical issue to be addressed.
In this study, we present the YOLO-TLA model, an improvement of YOLOv5 tailored to address the previously mentioned challenges. Key enhancements include the integration of an attention mechanism within the backbone network for better feature extraction and the fortification of multi-scale feature fusion in the neck network for more accurate detection of small objects. However, these improvements may lead to a rise in the number of parameters and computational demands. To address this, we propose and evaluate four lightweight strategies to find the optimal balance between detection performance and model complexity. The chosen strategy is then applied to YOLO-TLA to ensure an efficient yet effective model.
The main contributions of this paper are as follows:
  1. 1.
@@
  3. 3.
    Conducting a comparative study on the effectiveness of seven prominent attention mechanisms in the context of the YOLOv5s model.
  4. 4.
@@
II Related works
II-A Object detection
@@
R-CNN [14], developed by Girshick et al. , was a significant step, using a pre-trained CNN for feature extraction in various candidate regions and SVMs for classification. Despite its innovation, R-CNN suffered from redundant and complex processing, consuming extensive computational resources and storage. This led to the development of Fast R-CNN [15], which introduced the region proposal network (RPN) to replace the selective search approach. RPN generates candidate regions directly from the feature map, minimizing duplicated computations, and thereby enhancing model efficiency and speed. Moreover, Fast R-CNN employs a single feature extraction network for both classifying candidate boxes and performing bounding box regression, improving the use of features and detection accuracy. Advancing further, Ren et al. proposed Faster R-CNN [16], a landmark model in object detection. Faster R-CNN innovatively integrates region proposal generation with the feature extraction network, enabling comprehensive end-to-end training covering region proposal, feature extraction, classification, and bounding box regression. YOLO [17], proposed by Joseph et al. , approaches object detection by initially partitioning the input image into several grids. Each grid predicts potential bounding boxes for objects, spanning the entire image. These predictions are then refined through NMS to derive the final detection results. Liu et al. developed SSD [18], enhancing the VGG16 network architecture for multi-scale feature map detection. SSD leverages this enhanced backbone network to extract feature maps at various levels, each with distinct scales and features. Detection and classification are performed on these maps, and their predictions are aggregated to form the final detection outcome. YOLOv3 [19], proposed by Redmon et al. , marks a significant advancement by replacing its feature extraction network with DarkNet53, a network architecture known for its depth and superior feature extraction capabilities. YOLOv3 also leverages the feature pyramid network (FPNet) [20], using multi-scale downsampled feature maps for enhanced detection of small objects, thereby improving both inference speed and accuracy. On the other hand,Duan et al. introduced CenterNet [21], which eschews NMS and RPN, reducing accuracy degradation. CenterNet focuses on detecting object centers and calculating final bounding boxes with corner points and offset guidance. Tan et al. then proposed EfficientDet [22], which incorporates EfficientNet as its backbone network and introduces a bi-directional weighted feature pyramid network (BiFPN). The BiFPN allows for better feature representation through bidirectional cross-scale connectivity and efficient fusion, significantly enhancing detection speed while maintaining accuracy. Bochkovskiy et al. introduced YOLOv4 [23], a detection model that upgrades its backbone network to CSPDarknet53, an enhanced version of DarkNet53 featuring a CSP module. This enhancement results in a more modular network, simplifies the backbone network architecture, and augments feature extraction capabilities. YOLOv4 further evolves from YOLOv3 by adding a path aggregation network (PAN) architecture to its neck, drawing from the principles of PANet [24] to enhance multi-scale feature fusion. Ge et al. proposed YOLOx [10], building upon YOLOv5 while maintaining many of its design philosophy. YOLOv5 excels in single-stage object detection with its high accuracy and swift detection capabilities. YOLOx modifies YOLOv5 by implementing a decoupled detection head, separating the predictions of object class, bounding box, and intersection ratio into distinct branches. This decoupling has been found to not only enhance detection accuracy but also speed up the model convergence. On a different note, Dosovitskiy et al. pioneered the application of Transformer in computer vision with the Vision Transformer (ViT) [25], revolutionizing the research field of object detection. ViT adopts an encoder-decoder and self-attention mechanism to capture global image features, thus enabling completely end-to-end detection. However, this innovation comes at the cost of increased model parameters, raising training and deployment challenges. Concurrently, Carion et al. introduced DETR [26], the first to apply the Transformer architecture to object detection. DETR consists of a CNN-based backbone, an encoder, a decoder, and a feed forward network, forming a comprehensive system for detection. Despite its high parameter count and limited accuracy enhancements, DETR has been a cornerstone for subsequent advancements in Transformer-based object detection. Gold-YOLO [27], proposed by Wang et al. , represents a departure from the conventional YOLO model improvements. It introduces the gather-and-distribute mechanism to the neck network, significantly boosting the feature fusion capabilities and enabling more efficient information interaction. In a different approach, Chen et al. developed DiffusionDet [28], a pioneering effort in applying diffusion approaches to object detection. It redefines the detection task as a stepwise process of bounding box generation, utilizing decoupled training and iterative dynamic box evaluation. Zheng et al. introduced Focus-DETR [29], a lightweight modification of DETR. Focus-DETR addresses high complexity and parameter count that DETR faced by selectively processing object feature vectors, prioritizing only key objects. This selective approach can be tailored by adjusting the number of input vectors.
@@
Object detection methods based on CNNs typically utilize deep and wide backbone networks for feature extraction. They leverage multi-scale feature fusion [30] to capture extensive semantic information without neglecting geometric texture details, thereby enhancing the expressiveness of the detection feature map. However, these method come with a trade-off, as the numerous convolutional operations and stacked feature extraction networks significantly increase the model complexity and parameter count. On the other hand, Transformer-based object detection methods utilize self-attention mechanisms to model the interrelationships of different feature maps and globally contextualize them. These method are adept at assimilating information from multi-scale receptive fields. However, the complexity inherent in the Transformer architecture, particularly the interactions among several feature vectors, results in an increased number of parameters and higher computational load. Moreover, challenges such as object occlusion and background distraction persist in detection tasks. Therefore, current research primarily focuses on enhancing detection accuracy and robustness, achieving real-time processing, and reducing model complexity. This paper aims to bolster the detection of small objects by augmenting the feature extraction capabilities of the backbone network and the multi-scale feature fusion in the neck network. Simultaneously, we intend to decrease the overall parameters and computational demands through the application of lightweighting strategies.
II-B Attention mechanisms
@@
Figure 1: Overview network architecture of the proposed YOLO-TLA. In the figure, $k$, $s$, and $p$ indicate the convolutional kernel size, stride, and padding size, respectively.
III The proposed YOLO-TLA
III-A Motivation and baseline
In this study, we propose YOLO-TLA, an improved object detection model based on YOLOv5, with a focus on small object detection and reduced model complexity, as outlined in Fig.1. YOLOv5 presents five versions, named YOLOv5n, YOLOv5s, YOLOv5m, YOLOv5l, and YOLOv5x, in order of increasing size. Each version is designed with specific configurations to cater to different size needs. The model is structured into three primary parts: the backbone network, the neck network, and the head network. The backbone network is built on CSPDarknet53, consisting of standard convolutional layers with additional feature enhancement modules, tasked with extracting geometric texture features like the shape and color of objects. To enrich this basic information, the neck network, drawing inspiration from FPNet [20] and PANet [24], further combining feature maps from the backbone network with deeper semantic information. This combination results in feature maps rich in both semantic and geometric information. These enhanced feature maps are then fed into the head network, which performs the final detection and classification.
III-B Tiny object detection layer
In the MS COCO dataset [38], objects are categorized based on size: small objects are less than 32$\times$32 pixels, medium objects range from 32$\times$32 to 96$\times$ 96 pixels, and large objects exceed 96$\times$96 pixels. To enhance small object detection in our model, we have adjusted the feature map and anchor sizes. Utilizing the k-means clustering algorithm, we recalibrated the size range of preset anchor boxes and introduced a tiny object detection layer, leading to the creation of the YOLOv5s-Tiny model. Specifically, we upsample in the neck network of YOLOv5s to produce a 160$\times$160 feature map with 128 channels, which is then combined with the third layer output of the backbone network, matching in channels and size. This combined feature map, along with other detection layer outputs, is processed in the head network for classification and detection. Our neck network generates several feature maps at different scales, each corresponding to various anchor sizes. This enables the detection of smaller objects on larger feature maps and larger objects on smaller feature maps, enhancing object representation in the image. The correlation between feature map sizes and anchor sizes is detailed in Table I.
TABLE I: Anchor box sizes in different feature maps.
@@ III-B Tiny object detection layer
III-C Lightweight convolution module
@@
III-C1 C3 module revisited
The C3 module, a key component of YOLOv5, shown in Fig.2, comprises three standard convolutional layers, each with a 1$\times$1 kernel size and stride of 1, and includes several stacked BottleNeck modules. The architecture of this module varies in width and depth depending on the model size, controlled by predefined parameters. The C3 module incorporates a residual structure akin to BottleNeckCSP. It processes the input feature map in two ways: either through a dual-branch approach, where single-branch uses two standard convolutional layers and the other outputs the original feature map, subsequently concatenating the outputs, or by forgoing the residual path and directly outputting the feature map after standard convolutions. The BottleNeck module in C3 is noted for its robust feature extraction capabilities and its role in addressing gradient vanishing and explosion challenges.
@@ III-C3 C3CrossCovn module
In the MS COCO dataset, all images are RGB, so we use three image channels ($C=3$). To ensure a sufficiently large receptive field, we set $k$ to 3 and $s$ to 1. Under these conditions, it becomes apparent that standard convolution requires more computational effort and has approximately 1.5 times the number of parameters compared to CrossCovn. Although CrossCovn entails two striped-kernel convolution operations on a single feature map, it achieves finer feature extraction and richer feature information than standard convolution. This enhancement not only increases the detection accuracy but also significantly lowers computational demands and parameter count, making it an optimal solution for model lightweighting. The overview of the C3CrossCovn module is depicted in Fig.5.
@@
III-D Global attention mechanism
@@ III-E Detection pipeline
IV Experiments
IV-A Implement Details
In the experiments conducted for this study, the MS COCO dataset [38] was employed. Specifically, the COCO training set, comprising 118,287 images and 117,266 labels, and the COCO validation set, with 5,000 images and 4,952 labels, were used. Given the varying sizes of these images, they were uniformly resized to 640$\times$640 pixels for the experiments. The COCO training set includes a broad range of objects, totaling 80 object categories, which represent a comprehensive collection of common objects in daily life. This makes the MS COCO dataset widely applicable and valuable for research in computer vision. For training the model, the Adam optimizer is employed, starting with a learning rate of 0.001 and increasing to 0.01. To enhance the speed of parameter updates, we set the momentum to 0.937. Weight decay, crucial for regularization in training, is carefully balanced at 0.0005 to avoid overfitting or underfitting the model. The training regimen includes an initial warm-up phase covering the first 3 epochs, followed by an extensive training period spanning a total of 400 epochs. The model was implemented in Python, utilizing the PyTorch framework, and underwent training on a cloud server outfitted with four 16GB Tesla V100 GPUs. To expedite the experimental process while maintaining result accuracy, we employed different levels of data augmentation tailored to the varying model sizes.
IV-B Evaluation metrics
Several well-established metrics were chosen for evaluating the model performance in object detection tasks, including precision, recall, F1 score, and mean average precision (mAP), all of which serve to measure the efficacy of detection. For assessing the complexity of the model, metrics like the count of parameters and floating-point operations (FLOPs) were utilized. Precision and recall metrics are derived from a confusion matrix, which categorizes predictions into four types based on ground-truth and predicted labels. True positives (TP) occur when both the ground-truth and the prediction are positive. False negatives (FN) happen when the ground-truth is actually positive, but the prediction incorrectly indicates negative. False positives (FP) are cases where the ground-truth is actually negative, but the prediction is incorrectly positive. Finally, True negatives (TN) are instances where both the ground-truth and the prediction are negative. Precision and recall are calculated as follows,

### v2

@@ Abstract
1 Introduction
@@
Currently, there are two principal methodologies for implementing object detection models: the two-stage and the single-stage methods. The two-stage detection process starts with extracting image features using a convolutional neural network (CNN), followed by the generation of multiple candidate bounding boxes (or regions) on the feature map. Each box is then processed by additional convolutional layers for object classification and bounding box refinements. This includes identifying the object class and verifying its presence in the bounding box, and a regression phase to enhance the detection accuracy. The process concludes with non-maximum suppression (NMS), which filters out surplus bounding boxes to identify the most confident and accurate final detections. In contrast, the single-stage detection method directly predicts object class and their locations in the image, foregoing candidate bounding boxes generation. This method is generally faster and less computationally intensive than the two-stage method, making it ideal for scenarios requiring high real-time performance.
@@
Since its introduction, the YOLO series of object detection methods has garnered significant attention in the computer vision community, with several iterations enhancing its performance in both academic and industrial settings [9, 10, 11]. The detection model, YOLOv5, proposed by Ultralytics Inc., boasts improved accuracy and a more streamlined network architecture [12, 13, 14, 15]. Despite the advancements in YOLOv5, including its balance of speed and accuracy, it confronts specific challenges in industrial applications [16]. The model often struggles with accurately predicting the positions and classes of small or dense objects, negatively impacting detection performance. Furthermore, deploying these models in embedded systems, which typically have limited resources, necessitates a reduction in the number of parameters and computational demands. Consequently, optimizing the model to be more resource-efficient while preserving accuracy is a critical issue to be addressed. Ultralytics Inc. recognized these issues following the release of the YOLOv5 model and introduced an improved version, YOLOv8 [17, 18, 19, 20]. Both generations share similar inference processes and have the same backbone and neck network architectures. However, YOLOv8 introduces two major innovations compared to YOLOv5. Firstly, YOLOv8 employs a more efficient and lighter C2f module in the network architecture, corresponding to the C3 module in YOLOv5. The C2f retains the residual module and adds connections between different scale feature layers. Additionally, the split method is used to replace the convolution operation in the original residual branch. These improvements enable the C2f module to achieve richer gradient flows, enhance feature extraction capabilities, and reduce parameters. Secondly, YOLOv8 replaces the coupled architecture with the currently popular decoupled architecture in the head network and transitions the detection head from anchor-based to anchor-free. These modifications refine the detection process and improve the efficiency and accuracy.
In this paper, we present the YOLO-TLA object detection model based on the YOLOv5 model, which achieves higher accuracy and robustness while maintaining moderate model complexity in terms of parameters and computations. A distinctive improvement of YOLO-TLA is the tiny detection layer in the neck network, enhancing the multi-scale feature fusion process. This allows the model to better focus on small objects in the image. Additionally, YOLO-TLA employs a backbone network enhancement strategy that incorporates attention mechanisms at four locations in the backbone network to improve feature extraction capabilities. By comparing several mainstream attention mechanisms integrated with the YOLOv5s model and applying the best-performing global attention mechanism (GAM) [21] to the YOLO-TLA backbone network. Although these improvements positively impact detection capabilities, they inevitably increase the parameters and computational costs. Inspired by the C2f module design in the YOLOv8 model, we enhanced the feature extraction capabilities of the internal modules and simplified their complexity to improve model performance effectively. Specifically, we replace some C3 modules in the YOLOv5 network with lightweight feature extraction modules, C3CrossConv and C3Ghost, proposing four lightweight strategies and exploring their impact on model detection performance and model size. In summary, this study makes the following contributions:
@@
  • Comparing the impact of seven current mainstream attention mechanisms on the detection performance and complexity of the YOLOv5s model.
@@
  • Applying the aforementioned optimizations and improvements to the YOLOv5s model, proposing the final optimized model YOLO-TLAs.
  • Extending the design method of YOLO-TLA to the larger YOLOv5m model, proposing the YOLO-TLAm model.
The remainder of this paper is organized as follows. Existing object detection methods are introduced in Section 2. The following section, Section 3, proposes a small object detection model based on YOLOv5. In Section 4, experimental results and discussions demonstrate that the proposed method is efficient for small object detection. Finally, a brief conclusion is provided in Section 5.
2 Related works
2.1 Object detection
@@
R-CNN [22], developed by Girshick et al. , was a significant step, using a pre-trained CNN for feature extraction in various candidate regions and SVMs for classification. Despite its innovation, R-CNN suffered from redundant and complex processing, consuming extensive computational resources and storage. This led to the development of Fast R-CNN [23], which introduced the region proposal network (RPN) to replace the selective search approach. RPN generates candidate regions directly from the feature map, minimizing duplicated computations, and thereby enhancing model efficiency and speed. Moreover, Fast R-CNN employs a single feature extraction network for both classifying candidate boxes and performing bounding box regression, improving the use of features and detection accuracy. Advancing further, Ren et al. proposed Faster R-CNN [24], a landmark model in object detection. Faster R-CNN innovatively integrates region proposal generation with the feature extraction network, enabling comprehensive end-to-end training covering region proposal, feature extraction, classification, and bounding box regression. YOLO [25], proposed by Redmon et al. , approaches object detection by initially partitioning the input image into several grids. Each grid predicts potential bounding boxes for objects, spanning the entire image. These predictions are then refined through NMS to derive the final detection results. Liu et al. developed SSD [26], enhancing the VGG16 network architecture for multi-scale feature map detection. SSD leverages this enhanced backbone network to extract feature maps at various levels, each with distinct scales and features. Detection and classification are performed on these maps, and their predictions are aggregated to form the final detection outcome. YOLOv3 [27], proposed by Redmon et al. , marks a significant advancement by replacing its feature extraction network with DarkNet53, a network architecture known for its depth and superior feature extraction capabilities. YOLOv3 also leverages the feature pyramid network (FPNet) [28], using multi-scale downsampled feature maps for enhanced detection of small objects, thereby improving both inference speed and accuracy. On the other hand, Duan et al. introduced CenterNet [29], which eschews NMS and RPN, reducing accuracy degradation. CenterNet focuses on detecting object centers and calculating final bounding boxes with corner points and offset guidance. Tan et al. then proposed EfficientDet [30], which incorporates EfficientNet as its backbone network and introduces a bi-directional weighted feature pyramid network (BiFPN). The BiFPN allows for better feature representation through bidirectional cross-scale connectivity and efficient fusion, significantly enhancing detection speed while maintaining accuracy. Bochkovskiy et al. introduced YOLOv4 [9], a detection model that upgrades its backbone network to CSPDarknet53, an enhanced version of DarkNet53 featuring a CSP module. This enhancement results in a more modular network simplifies the backbone network architecture, and augments feature extraction capabilities. YOLOv4 further evolves from YOLOv3 by adding a path aggregation network (PAN) architecture to its neck, drawing from the principles of PANet [31] to enhance multi-scale feature fusion. Ge et al. proposed YOLOx [10], building upon YOLOv5 while maintaining much of its design philosophy. YOLOv5 excels in single-stage object detection with its high accuracy and swift detection capabilities. YOLOx modifies YOLOv5 by implementing a decoupled detection head, separating the predictions of object class, bounding box, and intersection ratio into distinct branches. This decoupling has been found to not only enhance detection accuracy but also speed up model convergence. On a different note, Dosovitskiy et al. pioneered the application of Transformer in computer vision with the Vision Transformer (ViT) [32], revolutionizing the research field of object detection. ViT adopts an encoder-decoder and self-attention mechanism to capture global image features, thus enabling complete end-to-end detection. However, this innovation comes at the cost of increased model parameters, raising training and deployment challenges. Concurrently, Carion et al. introduced DETR [33], the first to apply the Transformer architecture to object detection. DETR consists of a CNN-based backbone, an encoder, a decoder, and a feed forward network, forming a comprehensive system for detection. Despite its high parameter count and limited accuracy enhancements, DETR has been a cornerstone for subsequent advancements in Transformer-based object detection. Gold-YOLO [34], proposed by Wang et al. , represents a departure from the conventional YOLO model improvements. It introduces the gather-and-distribute mechanism to the neck network, significantly boosting the feature fusion capabilities and enabling more efficient information interaction. In a different approach, Chen et al. developed DiffusionDet [35], a pioneering effort in applying diffusion approaches to object detection. It redefines the detection task as a stepwise process of bounding box generation, utilizing decoupled training and iterative dynamic box evaluation. Zheng et al. introduced Focus-DETR [36], a lightweight modification of DETR. Focus-DETR addresses the high complexity and parameter count that DETR faces by selectively processing object feature vectors, prioritizing only key objects. This selective approach can be tailored by adjusting the number of input vectors.
@@
Object detection methods based on CNNs typically utilize deep and wide backbone networks for feature extraction. They leverage multi-scale feature fusion [37] to capture extensive semantic information without neglecting geometric texture details, thereby enhancing the expressiveness of the detection feature map. However, these methods come with a trade-off, as the numerous convolutional operations and stacked feature extraction networks significantly increase the model complexity and parameter count. On the other hand, Transformer-based object detection methods utilize self-attention mechanisms to model the interrelationships of different feature maps and globally contextualize them. These methods are adept at assimilating information from multi-scale receptive fields. However, the complexity inherent in the Transformer architecture, particularly the interactions among several feature vectors, results in an increased number of parameters and higher computational load. Moreover, challenges such as object occlusion and background distraction persist in detection tasks. Therefore, current research primarily focuses on enhancing detection accuracy and robustness, achieving real-time processing, and reducing model complexity. This paper aims to bolster the detection of small objects by augmenting the feature extraction capabilities of the backbone network and the multi-scale feature fusion in the neck network. Simultaneously, we intend to decrease the overall parameters and computational demands through the application of lightweight strategies.
2.2 Attention mechanisms
@@
3 Methodology
3.1 Motivation and baseline
Figure 1: Overview network architecture of the proposed YOLO-TLA. In the figure, $k$, $s$, and $p$ indicate the convolutional kernel size, stride, and padding size, respectively. The light green regions show our main improvements to the baseline method YOLOv5.
In this study, we propose YOLO-TLA, an improved object detection model based on YOLOv5, with a focus on small object detection and reduced model complexity, as outlined in Fig.1. YOLOv5 presents five versions, named YOLOv5n, YOLOv5s, YOLOv5m, YOLOv5l, and YOLOv5x, in order of increasing size. Each version is designed with specific configurations to cater to the needs of different sizes. The model is structured into three primary parts: the backbone network, the neck network, and the head network. The backbone network is built on CSPDarknet53, consisting of standard convolutional layers with additional feature enhancement modules, tasked with extracting geometric texture features like the shape and color of objects. To enrich this basic information, the neck network, drawing inspiration from FPNet [28] and PANet [31], further combines feature maps from the backbone network with deeper semantic information. This combination results in feature maps rich in both semantic and geometric information. These enhanced feature maps are then fed into the head network, which performs the final detection and classification.
3.2 Tiny object detection layer
In the MS COCO dataset [45], objects are categorized based on size: small objects are less than 32$\times$32 pixels, medium objects range from 32$\times$32 to 96$\times$ 96 pixels, and large objects exceed 96$\times$96 pixels. To enhance small object detection in our model, we have adjusted the feature map and anchor sizes. Utilizing the k-means clustering algorithm, we recalibrated the size range of preset anchor boxes and introduced a tiny object detection layer, leading to the creation of the YOLOv5s-Tiny model. Specifically, we upsample in the neck network of YOLOv5s to produce a 160$\times$160 feature map with 128 channels, which is then combined with the third layer output of the backbone network, matching in channels and size. This combined feature map, along with other detection layer outputs, is processed in the head network for classification and detection. Our neck network generates several feature maps at different scales, each corresponding to various anchor sizes. This enables the detection of smaller objects on larger feature maps and larger objects on smaller feature maps, enhancing object representation in the image. The correlation between feature map sizes and anchor sizes is detailed in Table 1.
Table 1: Anchor box sizes in different feature maps.
@@ 3.2 Tiny object detection layer
3.3 Lightweight convolution module
@@
3.3.1 C3 module revisited
The C3 module, a key component of YOLOv5, shown in Fig.2, comprises three standard convolutional layers, each with a 1$\times$1 kernel size and stride of 1, and includes several stacked BottleNeck modules. The architecture of this module varies in width and depth depending on the model size, controlled by predefined parameters. The C3 module incorporates a residual structure akin to BottleNeckCSP. It processes the input feature map in two ways: either through a dual-branch approach, where a single branch uses two standard convolutional layers and the other outputs the original feature map, subsequently concatenating the outputs, or by forgoing the residual path and directly outputting the feature map after standard convolutions. The BottleNeck module in C3 is noted for its robust feature extraction capabilities and its role in addressing gradient vanishing and explosion challenges.
@@ 3.3.3 C3CrossCovn module
In the MS COCO dataset, all images are RGB, so we use three image channels ($C=3$). To ensure a sufficiently large receptive field, we set $k$ to 3 and $s$ to 1. Under these conditions, it becomes apparent that standard convolution requires more computational effort and has approximately 1.5 times the number of parameters compared to CrossCovn. Although CrossCovn entails two striped-kernel convolution operations on a single feature map, it achieves finer feature extraction and richer feature information than standard convolution. This enhancement not only increases the detection accuracy but also significantly lowers computational demands and parameter count, making it an optimal solution for model lightweight. The overview of the C3CrossCovn module is depicted in Fig.5.
@@
3.4 Global attention mechanism
@@ 3.5 Detection pipeline
4 Experiments
4.1 Implement Details
In the experiments conducted for this study, the MS COCO dataset [45] was employed. Specifically, the COCO training set, comprising 118,287 images and 117,266 labels, and the COCO validation set, with 5,000 images and 4,952 labels, were used. Given the varying sizes of these images, they were uniformly resized to 640$\times$640 pixels for the experiments. The COCO training set includes a broad range of objects, totaling 80 object categories, which represent a comprehensive collection of common objects in daily life. This makes the MS COCO dataset widely applicable and valuable for research in computer vision. For training the model, the Adam optimizer is employed, starting with a learning rate of 0.001 and increasing to 0.01. To enhance the speed of parameter updates, we set the momentum to 0.937. Weight decay, crucial for regularization in training, is carefully balanced at 0.0005 to avoid overfitting or underfitting the model. The training regimen includes an initial warm-up phase covering the first 3 epochs, followed by an extensive training period spanning a total of 400 epochs. The model was implemented in Python, utilizing the PyTorch framework, and underwent training on a cloud server outfitted with four 16GB Tesla V100 GPUs. To expedite the experimental process while maintaining result accuracy, we employed different levels of data augmentation tailored to the varying model sizes. In addition, to ensure the accuracy and rigor of the experimental results and to exclude the interference of issues such as experimental hardware and software versions on the experimental results, all models mentioned in the experimental section of this paper, including the YOLOv5 model, were trained, validated, and tested using the same hardware platform and software environment.
4.2 Evaluation metrics
Several well-established metrics were chosen for evaluating the model performance in object detection tasks, including precision, recall, F1 score, and mean average precision (mAP), all of which serve to measure the efficacy of detection. Metrics like the count of parameters and floating-point operations (FLOPs) were utilized to assess the complexity of the model. Precision and recall metrics are derived from a confusion matrix, which categorizes predictions into four types based on ground-truth and predicted labels. True positives (TP) occur when both the ground-truth and the prediction are positive. False negatives (FN) happen when the ground-truth is actually positive, but the prediction incorrectly indicates negative. False positives (FP) are cases where the ground-truth is actually negative, but the prediction is incorrectly positive. Finally, True negatives (TN) are instances where both the ground-truth and the prediction are negative. Precision and recall are calculated as follows,